Calculate the amount of energy and KJ required to melt a 16.2 g aluminum pop can assume that the initial temperature of the aluminum it’s 20°C give your answer to two decimal places.

aditional info:

molar heat of fusion of Al=10.7 kj/mol

melting point of aluminum=660°c

heat capacity of Al=0.900j/g°c


1
Expert's answer
2021-04-19T05:11:15-0400

Ans:-

Mole"=\dfrac{16.2}{26.98154}g/mol=0.600 mol"

"\Delta H=(20\degree C\to660\degree C) +(melt)\\
\Delta H=(mc\Delta T)+(n\Delta Hfus)\\
\Delta H=(16.2g)(0.900J/g\degree C)(640\degree C)+(0.600 mol)(10.7kJ/mol)"

"\Delta H=9331.2J+6.42kJ\\
\Delta H=9.33kJ+6.42kJ\\
\Delta H=15.8kJ"